OpenStudy (anonymous):

Apply the quadratic formula to find the roots of the function and then graph it........ f(x) = x^2 - 4

OpenStudy (anonymous):

quadratic formula: (-b±sqrt(b^2-4ac))/(2a) where ax^2+bx+c

OpenStudy (anonymous):

so in this case, a=1, b=0, and c=-4 plugging it in, we get \[(-0±\sqrt{0^2-4(1)(-4)})/(2(1))\]

OpenStudy (anonymous):

so the first root is \[+\sqrt{16}/2=2\] and the second root is \[-\sqrt{16}/2=-2\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

I just really wish I understood any of that...lol............Oh well, it was worth a try...

OpenStudy (anonymous):

haha okay do you know the quadratic equation?

OpenStudy (anonymous):

formula*

OpenStudy (anonymous):

picture in your head a 43 year old woman who only managed to squeak through basic and advanced algebra, I am now teetering the line in the college algebra for the second time and it does not look good, I have been able to figure out some of this stuff but the quadratic might as well be in ancient script for all I understand of it

OpenStudy (anonymous):

I can handle the factoring, but the basic quadratic and the completing the square are next to impossible to understand unless you are a mathmatician, I'm a graphic designer....